# An International Agreement to Prevent the Premature Creation of Artificial Superintelligence

### v1

@@ 5 Why this plan in particular?
We recognize that restricting research is controversial. Unfortunately, we think it is necessary to prevent the premature development of ASI. This is primarily a descriptive statement, not a normative one: if you want to stop the development of ASI, the narrow set of research aimed at this goal must be stopped. In the current AI development paradigm, AI capabilities increase from a combination of scaling the number of operations used and having better AI algorithms and efficiencies; limiting only one of these factors would allow AI capabilities to improve through the other.
@@
In short, AI alignment is probably a difficult technical problem, and it is hard to be confident about solutions. Pausing for a substantial period gives humanity time to be careful in this domain rather than rushing. Rushing likely increases the chance of catastrophe and makes it more difficult to slow, even if there are obvious warning signs.
One reason why AI alignment is a difficult technical problem is that iteration becomes impossible at sufficiently high AI capability levels. Currently, AI systems are not smart enough to successfully take over. If AI capabilities increase sufficiently, AI systems (potentially acting in concert with one another) will have the ability to take over if they tried. In the latter regime, misalignment would be unrecoverably catastrophic (yudkowsky2025ifanyone). A core alignment difficulty is successfully generalizing alignment methods across this gap. If the future field of AI alignment believes it has solutions that will generalize across the gap, it has only one opportunity to test this theory.
@@
One particular objection to this paper’s plan is that a short pause may suffice. Some argue that a halt to AI development of just a couple years or a single decade would allow key safety and alignment challenges to be solved using AI labor. Automating AI alignment research was the goal of OpenAI’s now-disbanded Superalignment team (leike2023introducing), and is a key focus of Anthropic’s Alignment Science team (anthropic2025alignment). AI companies themselves have not publicly provided detailed plans for how they plan to automate alignment research, but other researchers have. In particular, greenblatt2024case aspires to use potentially misaligned AIs to automate AI safety and alignment research, thus making significant progress in a small amount of wall-clock time (i.e., in a year, AIs do 10-humanity-year equivalents of labor). If implemented by a sufficiently cautious actor, we think such a plan might work, but we don’t think one should be very confident in the plan’s success, and it would be an irresponsible gamble with immense stakes. As discussed, the state of the AI alignment field is nascent and has made little progress so far. There are various reasons to expect this plan to fail, or at least to not be confident in its success. This topic has been debated at length in internet forums, and for the sake of concision, we will only briefly outline a couple arguments:
@@
  • Alignment problems for ASI look like they are difficult problems. Early AGIs might simply not be able to solve these problems quickly. More capable AGIs would stand a better chance of solving these problems, but creating such AIs takes on more catastrophic misalignment risk. (wentworth2025case; lermen2025why)
@@
  • Early AGIs may not be aligned or controllable. This plan requires that early AGIs either be aligned themselves and thus trying to help solve alignment challenges, or that they are misaligned but we are still able to get useful labor out of them (as the AI Control research agenda aims to do (greenblatt2024case)). Each of these conditions seems prima facie unlikely, and it seems difficult to obtain justifiable confidence in at least one of them working.
@@
Simply monitoring newly produced chips is insufficient for our goals. If new chip production were perfectly monitored starting tomorrow, today’s existing chips could potentially be used to train dangerous AI systems. This isn’t a matter of ensuring that a large relative fraction of chips fall under the monitoring regime; even if existing chips only made up a small portion of total chips, there is still a large absolute number of them.
@@
A separate concern relating to the present plan is that it relies too heavily on remaining in the current AI development paradigm where significant computational resources are needed to improve AI capabilities. There are early signs that this may be changing, such as the inference-scaling regime started with OpenAI’s o-series of models (jaech2024openai). If implemented soon, this agreement would likely be sufficient, but the technical situation could deteriorate if action is deferred. Overall, our choice to go with this more permissive agreement—one that allows existing chips to keep existing and focuses primarily on preventing large training, is an attempt to balance costs and benefits, trying to preserve economically valuable inference on existing AI models without the risks from continued capabilities advancement.
@@ 6 Conclusion
We recognize that this proposal faces significant obstacles: it requires unprecedented international cooperation, restrictions on the use of AI chips, and limitations on scientific research. However, we believe these measures are proportionate to the risks involved and represent the most viable path to ensuring that ASI, when eventually developed, will be beneficial for humanity. The staged implementation approach detailed in our appendices provides a pathway forward even in the absence of immediate political consensus.
@@ Appendix A The Agreement
Below, we provide an annotated example draft language for the sort of agreement that could be implemented by major governments around the world, if they recognized the dangers from artificial superintelligence (ASI) and sought to prevent anyone from building ASI.
We are not policymakers and we are not well-versed in international law. We present this as an illustrative example of some potentially valuable provisions to have in view, using mechanisms tailored to the situation at hand and grounded in historical precedent.
@@
For each article, we’ve provided a commentary section explaining why we made key decisions, and a section discussing some relevant precedent.
@@
  17. 17.
    AI Technique Whitelist means the list of approved AI methods and techniques maintained by the Coalition Technical Body. Training runs between $10^{22}$ and $10^{24}$ FLOP may only use techniques on this list.
@@
Taiwan complicates our agreement concept, given its delicate geopolitical situation and its status as the producer of most of the world’s AI chips. Fortunately, precedent provides guidance: Though Taiwan is not a party to the NPT, it has stated on multiple occasions that it considers itself bound by the principles of the NPT. Taiwan allows the IAEA to conduct inspections and apply safeguards to its nuclear facilities through a trilateral agreement with the United States and the IAEA. A similar arrangement could be worked out with regards to this agreement.
@@
    Each Party agrees to ban and prohibit AI training above the following thresholds: Any training run exceeding $10^{24}$ FLOP or any post-training run exceeding $10^{23}$ FLOP. Each Party agrees to not conduct training runs above these thresholds, and to not permit any entity within its jurisdiction to conduct training runs above these thresholds.
@@
    Each Party shall report any training run between $10^{22}$ and $10^{24}$ FLOP to the CTB, prior to initiation. This applies for training runs conducted by the Party or any entity within its jurisdiction.
@@
      Training runs between $10^{22}$ and $10^{24}$ FLOP require explicit approval from inspectors representing the U.S. and the PRC, physically present at the training facility. Inspectors must provide written approval before training may commence. Either may withhold approval or require modifications. Inspectors shall verify that the training code uses only AI techniques approved on the AI Technique Whitelist maintained by the CTB.
@@
    Each Party, and entities within its jurisdiction, may conduct training runs of less than $10^{22}$ FLOP without oversight or approval from the CTB.
@@
    The CTB may authorize specific carveouts for activities such as safety evaluations, self-driving vehicles, medical technology, and other activities deemed safe by the Director-General, subject to the Executive Council’s veto power under Article III. These carveouts may allow for training runs larger than $10^{24}$ FLOP with CTB oversight, or a presumption of approval from the CTB for training runs between $10^{22}$ and $10^{24}$ FLOP.
@@
    The CTB creates and maintains an AI Technique Whitelist specifying allowed AI methods and techniques. The CTB may modify this Whitelist in accordance with Article III. Training runs between $10^{22}$ and $10^{24}$ FLOP may only employ techniques on this Whitelist.
@@
The hard prohibition at $10^{24}$ FLOP for training is slightly below that used to train models near the state of the art as of August 2025 (such as DeepSeek-V3, trained with $3\times{10^{24}}$ FLOP). We suggest this threshold because it is below the level at which we expect AIs to be dangerous (given current algorithms), and because it provides some breathing room and a buffer against algorithmic progress.
The prohibition of post-training over $10^{23}$ FLOP is meant to apply to the post-training of AIs created prior to the agreement entering into force. Many of these AIs will have been trained using more than $10^{24}$ FLOP; as of mid-2025, there are between 50 and 100 such models. Given that the weights of many such AIs will have been openly released, it is not feasible to prevent people from using them, but it is feasible to prevent large modifications to them via post-training.
A $10^{22}$ FLOP training run on 16 H100s would take around one week. This is computing-intensive enough that hobbyists would not accidentally break the threshold by training small and allowed models. AIs trained at the $10^{22}$ scale (with current algorithms) have so far appeared to be innocuous, though that situation would need to be monitored, as it could change as a result of unprevented AI research progress. Requiring reporting but not approval for training runs in the $10^{22}$ to $10^{24}$ FLOP band seems to us to strike a balance between allowing people to achieve some of the benefits from training AIs in a way that presently looks safe, while retaining the ability to prevent the creation larger, potentially dangerous AIs.
Monitoring would also allow the CTB to stay (somewhat) up to date with any algorithmic progress that happens in spite of the bans, and Article XIII provides for evaluations of models trained in this band. This would help to apprise the CTB of trends in AI development and adjust the FLOP thresholds however merited.
@@
Another reason to pursue consolidation is to limit breakout potential (breakout is when a party attempts to create ASI before anyone else can stop them). Consolidated assets are easier to surveil and strike in a crisis of confidence, and the anticipation of this improves the effectiveness of deterrence-by-denial.
@@
Chip centralization is an important first step to restricting the development of artificial superintelligence. Centralizing chips in declared facilities enables further monitoring for how these chips are being used, or verification that they are powered off (if they are not safe to use). Centralization would also make it easier for parties to destroy these chips, as might become necessary under Article XII, if a Party persists in violating the agreement.
@@
  • Monitoring chip clusters larger than 16 H100s works well with the training FLOP thresholds in Article IV. Training with 16 H100s (FP8 precision, 50 percent utilization — realistic but optimistic parameters) would take 7.3 days to get to $10^{22}$ FLOP, and 2 years to get to $10^{24}$ FLOP. Therefore, it would be feasible for people to use undeclared chips to reach the bottom threshold, but it would be somewhat impractical for them to get to the prohibited training threshold.
@@
    No Party sells or transfers non-AI advanced computer chips or non-AI advanced computer chip manufacturing equipment to non-Party States or entities outside a Party State except as authorized and tracked by the CTB.
@@
The AI chip supply chain is narrow and specialized, making it feasible to monitor production. The vast majority of AI chips are designed by NVIDIA. The most advanced logic chips (the main processor) used in AI chips are almost all fabricated by TSMC — accounting for around 90 percent of market share. Most AI chips are fabricated on versions of TSMC’s five-nanometer process node, a node likely only supported by two or three manufacturing plants. EUV lithography machines, a critical component in advanced logic chip fabrication, are made exclusively by ASML. High-bandwidth memory (HBM), another key component to AI chips, is dominated by two or three companies. This narrow and technical supply chain would be relatively easy to monitor and hard to clandestinely replicate.
@@
    Parties accept continuous on-site verification of total chip usage at declared CCCs. The methods used for verification will be determined and updated by the CTB, in accordance with the process described in Article III. These methods may include, but are not limited to:
@@
      Measurements of power, thermal, and networking characteristics (i.e. off-chip monitoring mechanisms)
@@
This article creates a positive incentive to join the agreement: A country may continue using AI chips as long as supervision can verify that their use does not put the world at risk. Given the goal of preventing large-scale AI training, there are two main approaches: Ensure nobody has the necessary hardware (i.e., that AI chips do not exist), or ensure that the hardware is not used in the development of superintelligence (i.e. via monitoring). Monitoring is what permits the continued safe use of AI chips. This is conceptually analogous to IAEA Safeguards: In order for a non-nuclear weapon country to be permitted nuclear materials and facilities, it is necessary for the IAEA to inspect and ensure the use is only for peaceful purposes.
@@
    For the purpose of preventing the development of artificial superintelligence, this Agreement restricts only research that would materially advance toward ASI or undermine verification of compliance with this Agreement. Research focused on specific applications (such as medical diagnosis, scientific discovery, or industrial automation) that does not advance general cognitive capabilities toward ASI levels is not restricted. Restricted Research includes:
@@
      Research into computer artificial intelligence paradigms beyond machine learning
@@
Unlike other types of government classification, Restricted Data can be created (deliberately or accidentally) by the private sector, a matter of unresolved constitutionality [^13] The 1979 case of United States v. The Progressive, in which a newspaper intended to reveal the “secret” of the hydrogen bomb, might have given the U.S. Supreme Court an opportunity to rule on whether the “born secret” doctrine violates the First Amendment’s protections on speech, if the government hadn’t dropped the case as moot. that highlights the need for a regulatory arm authorized and capable of making everyday decisions about the exact boundaries of Restricted Data. The National Nuclear Security Administration (NNSA) does this for nuclear secrets in the U.S. Under our Article VIII, paragraph 5, the Research Controls division would take on this role for restricted AI research. It would also fill other NNSA-analogous functions, outlined in our Article IX, by (1) maintaining relationships with researchers and and organizations working on projects that approach the classification threshold, and (2) establishing secure infrastructure for reporting and containment of inadvertent discoveries.
@@
Some research must be banned to prevent AI capabilities from advancing, even when holding the amount of training FLOP constant. This ban would need to cover all research that might make AIs more efficient to train or that might increase the capabilities of AIs, often referred to as “algorithmic progress.” In current paradigms, this includes advances in the algorithms used in pre-training, post-training, and inference. As paradigms change, these distinctions may become less clear and new categories may arise. [^16] For instance, the development of AlphaGo [silver2016mastering] — a state-of-the-art AI in 2016 — doesn’t fit cleanly into the modern “pre-training, post-training, inference” paradigm. For this reason, the agreement uses the terms “development, deployment, or use.”
@@
        Monitoring of the employment status and whereabouts of researchers who have previously worked in Restricted Research topics, or are presently working in adjacent areas.
@@
Precedent for “monitoring of the employment status and whereabouts of researchers” in high-risk fields, as we suggest in paragraph 2, subparagraph (a)(ii), can be found in the International Science and Technology Center (ISTC). [^17] The International Science and Technology center grew out of the 1991 Nunn-Lugar Cooperative Threat Reduction program, a U.S. initiative to secure and dismantle WMDs and their associated infrastructure in former Soviet states. Established in 1994, the ISTC was specifically created to reduce nuclear proliferation risks by keeping Soviet nuclear researchers gainfully employed in peaceful activities and connected to the international scientific community. The ISTC also shows the potential of incentives as a complement to penalties for keeping technical experts (who may find themselves unemployed as a result of this agreement) from engaging in Restricted Research.
To the extent that penalties may need to be severe to provide the deterrence indicated in our Article IX, paragraph 1, subparagraph (b), a template may be found with the Enforcement chapter (18) of the 1946 Atomic Energy Act, under which the unauthorized sharing of Restricted Data can be punished by death or imprisonment if the disclosures were made with treasonous intent. [^18] Parties to our agreement may wish to explore expanding the concept of crimes against humanity (codified in the 1988 Rome Statute of the International Criminal Court) to cases where a researcher deliberately seeks to develop ASI at the expense of the people of Earth.
@@
The Research Controls division might look to existing practices by the IAEA when developing inspection protocols. Under the framework of the Model Additional Protocol approved in 1997 by the IAEA Board of Governors, states that have made comprehensive safeguard agreements202020144 States, as of June 2025 allow complementary access inspections that look for undeclared nuclear material. As part of such visits, inspectors may interview operators, analogous to our proposal in paragraph 2, subparagraph (a)(i).

### v2

@@ 5 Why this plan in particular?
We recognize that restricting research is controversial and normally a bad idea. Unfortunately, we think it is necessary to prevent the premature development of ASI. This is primarily a descriptive statement, not a normative one: to prevent the development of ASI, the narrow set of research aimed at this goal must be stopped. In the current paradigm, AI capabilities increase from a combination of scaling the number of operations used and having better AI algorithms; limiting only one of these factors would allow AI capabilities to improve through the other.
@@
In short, AI alignment is probably a difficult technical problem, and it is hard to be confident about solutions. Pausing for a substantial period gives humanity time to be careful in this domain rather than rushing. Pausing for a shorter amount of time (e.g., 5 years) might reduce risk substantially compared to the current race, but it also might not be enough. In general, world leaders should weigh the likelihood and consequence of different risks and benefits against each other for different lengths of a pause.
Section 2 discusses some of the reasons why the AI alignment problem may be difficult. Generally, experts vary in their estimates of the difficulty of this problem and the likelihood of catastrophe, with some expecting the problem to be very hard (grace2025thousands; controlai_quotes_2025; pdoom2025wikipedia). Given this uncertainty about how difficult this problem is, we should prepare to pause for a long time, in case more effort is needed. Our agreement would allow for a long halt, even if world leaders later came to believe a shorter one was acceptable.
We also contend that there are other problems which need to be addressed during a halt even if one presumes that alignment can be quickly solved, and these problems are also of an uncertain difficulty. These include risks of power concentration, human misuse of AIs, mass-unemployment, and many more. World leaders will likely want at least years to understand and address these problems. The international agreement proposed in this paper is primarily motivated by risks from AI misalignment, but there are numerous other risks that it would also help reduce.
What about automating AI alignment research?
@@
One objection to this paper’s plan is that a short pause may suffice because AI alignment research could be automated. Automating AI alignment research was the goal of OpenAI’s now-disbanded Superalignment team (leike2023introducing), and is a key focus of Anthropic’s Alignment Science team (anthropic2025alignment). AI companies themselves have not publicly provided detailed plans for how they plan to automate alignment research. Assuming they had such a plan, and assuming they acted with much more caution than the AI field of today, this plan might work. But we don’t think one should be very confident in the plan’s success, and it would be an irresponsible gamble with immense stakes. As discussed, the state of the AI alignment field is nascent and has made little progress so far. There are various reasons to expect this plan to fail, or at least to not be confident in its success. This topic has been debated at length in internet forums, and for the sake of concision, we will only briefly outline a couple arguments:
@@
  • Alignment problems for ASI appear difficult. Early AGIs may not be able to solve these problems quickly. More capable AGIs would have a better chance of solving these problems, but creating such AIs takes on more catastrophic misalignment risk. (wentworth2025case; lermen2025why)
@@
  • Early AGIs may not be aligned or controllable. This plan requires that early AGIs either be aligned themselves and thus trying to help solve alignment challenges, or that they are misaligned but we are still able to get useful labor out of them (as the AI Control research agenda aims to do greenblatt_ai_2024). Each of these conditions seems prima facie unlikely, and it seems difficult to obtain justifiable confidence in at least one of them working.
@@
Simply monitoring newly produced chips is insufficient for our goals due to the existing chip stock. One might argue that new chips will comprise a much larger fraction of the world’s computing power than existing chips. However, risk derives from the absolute number of unmonitored chips, not just their proportion in the total chip stock.
@@
Won’t this plan enable authoritarianism?
This agreement does include elements that could be misused for authoritarian abuse, such as centralizing AI infrastructure, monitoring training runs, and tracking researchers. However, the default trajectory carries far greater authoritarianism risk. ASI may be highly power-concentrating: if someone manages to develop an aligned superintelligence, it may grant them unrivalled economic, military, or technological power. Under current conditions, ASI will likely be built by a private company (potentially in partnership with government agencies), with minimal public oversight. In this situation, control over the most powerful technology would rest with perhaps 1-10 individuals. This would be an unprecedented concentration of power with few mechanisms to prevent permanent authoritarian lock-in. By contrast, this agreement prevents the creation of ASI, allowing time to establish the required governance structures.
The agreement’s verification mechanisms (e.g., tracking chips, monitoring training runs, limiting research) would likely involve a combination of existing and new authorities. As a few examples of related precedent, the U.S. government already conducts extensive surveillance (including against U.S. persons), maintains export controls on sensitive technologies, inspects nuclear facilities, and restricts the publication of nuclear research. Where domestic implementation expands the government’s power, there are likely safeguards that could be implemented to prevent misuse. For instance, multi-lateral controls, narrow scoping of authority, and transparency measures would reduce these risks.
@@
Another concern relating to the plan is that it relies too heavily on remaining in the current AI development paradigm where significant computational resources are needed to improve AI capabilities. There are early signs that this may be changing, such as the inference-scaling regime started with OpenAI’s o-series of models (jaech2024openai). If implemented soon, this agreement would likely be sufficient, but the technical situation could deteriorate if action is deferred. Overall, our choice to go with this more permissive agreement—one that allows existing chips to keep existing and focuses primarily on preventing large training—is an attempt to balance costs and benefits, trying to preserve the safe use of existing AI models without the risks from continued capabilities advancement.
@@ 6 Conclusion
We recognize that this proposal faces significant obstacles: it requires unprecedented international cooperation, restrictions on the use of AI chips, and limitations on scientific research. However, we believe these measures are proportionate to the risks and represent the most viable path to ensuring that ASI, when eventually developed, will be beneficial for humanity. The staged implementation in the appendices provides a pathway forward even in the absence of immediate political consensus.
@@
Acknowledgments
We are thankful to the following people and numerous others for feedback on earlier versions of this work. They do not necessarily endorse the views expressed here.
Nate Soares, Mitch Howe, Joseph Rogero, Robi Rahman, Rob Bensinger, Malo Bourgon, Daniel Kokotajlo, Thomas Larsen.
LLMs were used to help with various parts of this report, including background research, text drafting, and text refinement. All final text was reviewed by a human and the vast majority was written by a human.
@@ Appendix A The Agreement
Below, we provide an annotated example draft language for the sort of agreement that could be implemented by major governments around the world, if they recognized the dangers from artificial superintelligence (ASI) and sought to prevent anyone from building ASI. We present this as an illustrative example of some potentially valuable provisions to have in view, using mechanisms tailored to the situation at hand and grounded in historical precedent.
@@
For each article, we’ve provided a commentary section explaining why we made key decisions, and a section discussing some relevant precedent. It’s important to note that precedent is valuable because it demonstrates that the plan can be effectuated, not that it will be effective. The existence of prior agreements shows that similar mechanisms are politically and practically feasible, but does not guarantee they will achieve their intended outcomes in this domain. By contrast, our reasons for believing this particular agreement would be effective are discussed in various places, including the commentary sections, and are often AI-specific.
@@
    Strict Threshold is the amount of training computation (measured in FLOP) above which training runs are prohibited. It is set at $10^{24}$ FLOP.
  9. 9.
    Strict Post-training Threshold is the amount of training computation (measured in FLOP) above which post-training runs (e.g., of models trained before the agreement) are prohibited. It is set at $10^{23}$ FLOP.
  10. 10.
    Monitored Threshold is the amount of training computation (measured in FLOP) above which training runs are subject to monitoring by the international authority. It is set at $10^{22}$ FLOP.
  11. 11.
@@
  20. 20.
    AI Technique Whitelist means the list of approved AI methods and techniques maintained by the Coalition Technical Body. Training runs above the Monitored Threshold may only use techniques on this list.
@@
Taiwan complicates our agreement concept, given its delicate geopolitical situation and its status as the producer of most of the world’s AI chips. Fortunately, precedent provides guidance: Though Taiwan is not a party to the NPT, it has stated on multiple occasions that it considers itself bound by the principles of the NPT. Taiwan allows the IAEA to conduct inspections and apply safeguards to its nuclear facilities through a trilateral agreement with the United States and the IAEA. A similar arrangement could be worked out with regard to this agreement.
@@
    Each Party agrees to ban and prohibit AI training above the following thresholds: Any training run exceeding the Strict Threshold or any post-training run exceeding the Strict Post-training Threshold. Each Party agrees to not conduct training runs above these thresholds, and to not permit any entity within its jurisdiction to conduct training runs above these thresholds.
@@
    Each Party shall report any training run above the Monitored Threshold to the CTB, prior to initiation. This applies for training runs conducted by the Party or any entity within its jurisdiction.
@@
      Training runs above the Monitored Threshold require explicit approval from inspectors representing the U.S. and the PRC, physically present at the training facility. Inspectors must provide written approval before training may commence. Either may withhold approval or require modifications. Inspectors shall verify that the training code uses only AI techniques approved on the AI Technique Whitelist maintained by the CTB.
@@
    Each Party, and entities within its jurisdiction, may conduct training runs of less FLOP than the Monitored Threshold without oversight or approval from the CTB.
@@
    The CTB may authorize specific carveouts for activities such as safety evaluations, self-driving vehicles, medical technology, and other activities deemed safe by the Director-General, subject to the Executive Council’s veto power under Article III. These carveouts may allow for training runs larger than the Strict Threshold with CTB oversight, or a presumption of approval from the CTB for training runs between the Monitored Threshold and Strict Threshold.
@@
    The CTB creates and maintains an AI Technique Whitelist specifying allowed AI methods and techniques. The CTB may modify this Whitelist in accordance with Article III. Training runs above the Monitored Threshold may only employ techniques on this Whitelist.
@@
The hard prohibition at the Strict Threshold ($10^{24}$ FLOP) for training is slightly below that used to train models near the state of the art as of August 2025 (such as DeepSeek-V3, trained with $3\times{10^{24}}$ FLOP). We suggest this threshold because it is below the level at which we expect AIs to be dangerous (given current algorithms), and because it provides some breathing room and a buffer against algorithmic progress.
The prohibition of post-training over $10^{23}$ is meant to apply to the post-training of AIs created prior to the agreement entering into force. Many of these AIs will have been trained using more than $10^{24}$ FLOP; as of mid-2025, there are between 50 and 100 such models. Given that the weights of many such AIs will have been openly released, it is not feasible to prevent people from using them, but it is feasible to prevent large modifications to them via post-training.
A $10^{22}$ FLOP (Monitored Threshold) training run on 16 H100s would take around one week. This is computing-intensive enough that hobbyists would not accidentally break the threshold by training small and allowed models. AIs trained at the $10^{22}$ scale (with current algorithms) have so far appeared to be innocuous, though that situation would need to be monitored, as it could change as a result of unprevented AI research progress.
Monitoring of medium-scale training would allow the CTB to stay (somewhat) up to date with any algorithmic progress that happens in spite of the bans, and Article XIII provides for evaluations of models trained in this band. This would help to apprise the CTB of trends in AI development and adjust the FLOP thresholds however merited.
@@
Another reason to pursue consolidation is to limit breakout potential (breakout is when a party attempts to create ASI before anyone else can stop them). Consolidated assets are easier to surveil and disable if necessary. The anticipation of a credible disabling action improves the effectiveness of deterrence-by-denial.
@@
Chip centralization is an important first step to restricting the development of artificial superintelligence. Centralizing chips in declared facilities enables further monitoring for how these chips are being used, or verification that they are powered off (if they are not safe to use). Centralization would also make it easier for parties to disable these chips, as might become necessary under Article XII, if a Party persists in violating the agreement.
@@
  • Monitoring chip clusters larger than 16 H100s works well with the training FLOP thresholds in Article IV. Training with 16 H100s (FP8 precision, 50 percent utilization — realistic but optimistic parameters) would take 7.3 days to get to $10^{22}$ FLOP (the Monitored Threshold), and 2 years to get to $10^{24}$ FLOP (the Strict Threshold). Therefore, it would be feasible for people to use undeclared chips to reach the bottom threshold, but it would be somewhat impractical for them to get to the prohibited training threshold.
@@
    No Party sells or transfers non-AI advanced computer chips or non-AI advanced computer chip manufacturing equipment to non‑Party States or entities outside a Party State except as authorized and tracked by the CTB.
@@
The AI chip supply chain is narrow and specialized, making it feasible to monitor production. The vast majority of AI chips are designed by NVIDIA. The most advanced logic chips (the main processor) used in AI chips are almost all fabricated by TSMC — accounting for around 90 percent of market share. Most AI chips are fabricated on versions of TSMC’s five-nanometer process node, a node likely only supported by two or three manufacturing plants. EUV lithography machines, a critical component in advanced logic chip fabrication, are made exclusively by ASML. High-bandwidth memory (HBM), another key component to AI chips, is dominated by two or three companies. This narrow and technical supply chain would be relatively easy to monitor and hard to clandestinely replicate. We don’t want to overstate things too much—for example, China has an emerging domestic supply chain that produces some notable AI chips—but even with various caveats like this, monitoring existing chip production seems quite feasible.
@@
    Parties accept continuous on‑site verification of total chip usage at declared CCCs. The methods used for verification will be determined and updated by the CTB, in accordance with the process described in Article III. These methods may include, but are not limited to:
@@
      Measurements of power, thermal, and networking characteristics (i.e., off-chip monitoring mechanisms)
@@
This article creates a positive incentive to join the agreement: A country may continue using AI chips as long as supervision can verify that their use does not put the world at risk. Given the goal of preventing large-scale AI training, there are two main approaches: Ensure nobody has the necessary hardware (i.e., that AI chips do not exist), or ensure that the hardware is not used in the development of superintelligence (i.e., via monitoring). Monitoring is what permits the continued safe use of AI chips. This is conceptually analogous to IAEA Safeguards: In order for a non-nuclear weapon country to be permitted nuclear materials and facilities, it is necessary for the IAEA to inspect and ensure the use is only for peaceful purposes.
@@
    For the purpose of preventing the development of artificial superintelligence, this Agreement restricts only research that would materially advance toward ASI or undermine verification of compliance with this Agreement. This includes research in the field of machine learning and research in other artificial intelligence paradigms. Research focused on specific applications (such as medical diagnosis, scientific discovery, or industrial automation) that does not advance general cognitive capabilities toward ASI levels is not restricted. Restricted Research includes:
@@
Unlike other types of government classification, Restricted Data can be created (deliberately or accidentally) by the private sector, a matter of unresolved constitutionality [^13] The 1979 case of United States v. The Progressive, in which a newspaper intended to reveal the “secret” of the hydrogen bomb, might have given the U.S. Supreme Court an opportunity to rule on whether the “born secret” doctrine violates the First Amendment’s protections on speech, if the government hadn’t dropped the case as moot. that highlights the need for a regulatory arm authorized and capable of making everyday decisions about the exact boundaries of Restricted Data. The National Nuclear Security Administration (NNSA) does this for nuclear secrets in the U.S. Under our Article VIII, paragraph 5, the Research Controls division would take on this role for restricted AI research. It would also fill other NNSA-analogous functions, outlined in our Article IX, by (1) maintaining relationships with researchers and organizations working on projects that approach the classification threshold, and (2) establishing secure infrastructure for reporting and containment of inadvertent discoveries.
@@
The 1975 Asilomar Conference on Recombinant DNA is evidence that scientists themselves may voluntarily agree to research prohibitions, especially when faced with novel threats they are not sure they can contain. This conference resulted in voluntary guidelines around recombinant DNA research. These guidelines included prohibitions on certain especially dangerous experiments, such as cloning recombinant DNA from highly pathogenic organisms and DNA containing toxin genes.
@@
Some research must be banned to prevent AI capabilities from advancing, even when holding the amount of training FLOP constant. This ban would need to cover all research that might make AIs more efficient to train or that might increase the capabilities of AIs, often referred to as “algorithmic progress.” In current paradigms, this includes advances in the algorithms used in pre-training, post-training, and inference. The AI development paradigm might shift and other AI development paradigms might catch up to machine learning. Therefore, the agreement should not constrain itself to only research in machine learning, even if this is the most urgent priority today. The agreement keeps the door open to potentially restrict research in other AI paradigms (e.g., connectomics, brain-inspired AI, fast genetic algorithms, GOFAI), if research in those paradigms seems likely to lead to ASI.
@@
        Monitoring of the employment status of researchers who have previously worked in Restricted Research topics, or are presently working in adjacent areas.
@@
Precedent for “monitoring of the employment status of researchers” in high-risk fields, as we suggest in paragraph 2, subparagraph (a)(ii), can be found in the International Science and Technology Center (ISTC). [^16] The International Science and Technology center grew out of the 1991 Nunn-Lugar Cooperative Threat Reduction program, a U.S. initiative to secure and dismantle WMDs and their associated infrastructure in former Soviet states. Established in 1994, the ISTC was specifically created to reduce nuclear proliferation risks by keeping Soviet nuclear researchers gainfully employed in peaceful activities and connected to the international scientific community. The ISTC also shows the potential of incentives as a complement to penalties for keeping technical experts (who may find themselves unemployed as a result of this agreement) from engaging in Restricted Research.
Penalties may need to be severe to provide the deterrence indicated in our Article IX, paragraph 1, subparagraph (b). While it may not be directly applicable here, there is precedent for severely punishing the sharing of sensitive data in some contexts. The Enforcement chapter (18) of the 1946 Atomic Energy Act includes fines and imprisonment for unauthorized disclosures. More severe punishments are sometimes levied for other information-disclosure crimes, namely treason. [^17] Parties to our agreement may wish to explore expanding the concept of crimes against humanity (codified in the 1988 Rome Statute of the International Criminal Court) to cases where a researcher deliberately seeks to develop ASI at the expense of the people of Earth.
@@
The Research Controls division might look to existing practices by the IAEA when developing inspection protocols. Under the framework of the Model Additional Protocol approved in 1997 by the IAEA Board of Governors, states that have made comprehensive safeguard agreements191919144 States, as of June 2025 allow complementary access inspections that look for undeclared nuclear material. As part of such visits, inspectors may interview operators, analogous to our proposal in paragraph 2, subparagraph (a)(i).